# Untestability of Average Slutsky Symmetry

### v1

@@ Abstract
Slutsky symmetry and negative semidefiniteness are necessary and sufficient conditions for the rationality of demand functions. While the empirical implications of Slutsky negative semidefiniteness in repeated cross-sectional demand data are well understood, the empirical content of Slutsky symmetry remains largely unexplored. This paper takes an important first step toward addressing this gap. We demonstrate that the average Slutsky matrix is not identified and that its identified set always contains a symmetric matrix. A key implication of our findings is that the symmetry of the average Slutsky matrix is untestable, and consequently, individual Slutsky symmetry cannot be tested using the average Slutsky matrix.
Keywordsβ Slutsky symmetry, rationality, cross-sectional demand, continuity equation
@@ 1 Introduction
Rationality is a central concept in economic theory, serving as a fundamental assumption in the analysis of consumer behavior. It assumes that consumers make decisions consistent with maximizing utility, a principle that supports many economic models and empirical research. Testing this assumption is crucial for validating theoretical models and understanding consumer decision-making.
When individual demand functions are available, the Hurwicz-Uzawa theorem ([hurwicz1971integrability]) provides a complete characterization of rationality. To state the theorem, let be a -dimensional price vector and be an income, both of which are relative to the price of the numeraire. Let be the -dimensional vector of quantities demanded when the price is and the income is We say that a demand function is rational if it maximizes a utility function subject to the budget constraint. According to the Hurwicz-Uzawa theorem, a demand function is rational if and only if its Slutsky matrix, which is defined as a matrix
@@
In many real-world applications, individual demand functions are not observed and only cross-sectional data is available. To investigate the population rationality, researchers have explored the rationalizability of the average demand function conditional on observable characteristics, such as price and income, treating it as if it is generated by an individual representing the population. See, for example, [lewbel1995consistent], [lewbel2001demand], [haag2009testing], and [hoderlein2011many]. However, this approach involves some disadvantages. For instance, as Theorem 1 of [lewbel2001demand] demonstrates, the rationality of the average demand function is unrelated to the rationality of individual demand functions: it is possible for all individuals in a population to be rational while the average demand is not, and vice versa. Moreover, it is often reported that cross-sectional mean regressions fail to explain the variation of demand adequately because of the unobserved preference heterogeneity. ([hoderlein2011many], [hausman2016individual].) This fact indicates the need to investigate all the information about the heterogeneity contained in the data.
@@
Testing rationality for cross-sectional demand data is a harder problem than that for individual demands because individual Slutsky matrices are not observed. This difficulty raises a fundamental question: can we infer the rationality of a population from such data? More specifically, is it possible to construct statistical tests with nontrivial power to detect consumer irrationality using only distributional data?
@@
Recent research has made progress in addressing this question. [hausman2016individual] provide a necessary and sufficient condition for cross-sectional demand distributions to be rationalizable when only two goods are present. For cases involving more than two goods, [dette2016testing] provide the empirical content of the negative semidefiniteness of the Slutsky matrix. Specifically, they show that the quadratic form of the average Slutsky matrix conditional on price and income, where is the random demand function representing the population, is identified. If an estimate of the quadratic form is significantly positive, then it means that the observed data is likely to be inconsistent with a population having an negative semidefinite Slutsky matrixβlet alone rationality. These studies demonstrate the feasibility of testing rationality by focusing on the negative semidefiniteness of the Slutsky matrix. However, these methods neglect the symmetry condition, leading to statistical tests that are overly conservative and fail to fully leverage the implications of rationality. The symmetry condition can be more critical to the rationality than the negative semidefiniteness, given that the former is highly sensitive to small perturbations while the latter is not. If Slutsky symmetry is testable, one should be able to construct a much more powerful test for rationality by testing both conditions.
@@
Despite its theoretical importance, the empirical content of the Slutsky symmetry condition remains poorly understood. This paper provides a first step toward elucidating the implications of individual Slutsky symmetry for cross-sectional demand distributions. We demonstrate that the average Slutsky matrix is not identified unlike its quadratic form, which is identified as shown by [dette2016testing]. Moreover, its identified set always contains a symmetric matrix. To prove these results, we construct a stochastic demand system that is observationally equivalent to the true demand system while ensuring that the average Slutsky matrix is symmetric.
@@
An immediate consequence of our result is the fundamental untestability of individual Slutsky symmetry via the average Slutsky matrix. This fact indicates that the Slutsky symmetric condition is totally distinct to the negative semidefiniteness condition, which is testable based on the quadratic form of the average Slutsky matrix ([dette2016testing]). Note also that it does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. It remains an open question whether individual symmetry has observable implications through nonlinear statistics, a topic that is not explored in this paper.
The key theoretical contribution of this paper is the introduction of the continuity equation as a tool to construct a random demand function that satisfies given marginal distributions and adheres to conditions on the Slutsky matrix. The continuity equation originates from fluid dynamics, where it describes the time evolution of fluid density. It is expressed as:
where is the divergence operator defined as for a vector-valued function is the fluid density, and is the flow velocity vector field at time The first term of the LHS represents the change in fluid volume within the system, while the second term accounts for the net flow difference (inflow minus outflow). This equation describes situations where all changes in fluid volume are due to particle transport, with no external sources or sinks present. In the context of demand analysis, heterogeneous individuals are conceptualized as a fluid, with price and income interpreted as βmultidimensional time.β The continuity equation then captures how demand distributions evolve as price and income change. The velocity field corresponds to the derivatives of demand with respect to price and income, which are components of the Slutsky matrix. Our approach to constructing a desired random demand function involves finding a βvelocityβ field that obeys the continuity equation while satisfying the rationality condition.
@@
Related literature. [hurwicz1971integrability] investigate the integrability of individual demand functions and give a necessary and sufficient condition for rationality based on the Slutsky matrix. [lewbel2001demand] considers a population that is heterogeneous in preference and provides conditions for the average demand function to be rational. [haag2009testing] and [hoderlein2011many] among others consider the estimation and inference on the average demand function under the rationality shape restriction.
@@
For cross-sectional demand distributions, [hausman2016individual] give a necessary and sufficient condition for observed datasets to be consistent with a rational demand system for the cases when there are only two goods. When more than two goods are present, the characterization of rationalizability is largely open. [dette2016testing] provide a necessary condition that the data needs to satisfy for it to be rationalizable focusing on the negative semidefiniteness of the Slutsky matrix and propose a statistical testing for rationality. [maes2024beyond] construct a testing procedure of rationality based on higher order moments of demand distributions. They also observe that the average Slutsky matrix is not identified from those moments.
@@
This paper is also related to the literature on random utility models. [mcfadden1990stochastic] and [mcfadden2005revealed] show that the Axiom of Revealed Stochastic Preference characterizes rationalizability of stochastic choice functions defined on a finite number of choice sets. In a similar setup, [kitamura2018nonparametric] constructed a statistical test for the axiom.
In the context of discrete choice, [bhattacharya2025integrability] recently gives the complete characterization of the rationalizability of demand distributions.
2 Setup and Main Result
@@
For we consider an economy with goods. Relative to the first good, their prices are encoded into a price vector where Let be the relative income. For given price and income a consumer demands Notice that we assume the homogeneity of demand functions at this point. We also assume Warlasβ law, i.e., the demand for the numeraire is and consequently, is assumed.
@@
In the real world, consumerβs preference heterogeneity is present. In this sense, consumerβs demand is stochastic from the perspective of an econometrician. Let be a random individual demand function drawn from a probability distribution on The econometrician is assumed to observe cross-sectional demand distributions, that is, (s)he observes the distribution of for each but no joint distribution of demands across different price-income levels is available. Regarding as a stochastic process indexed by we often call the (one-dimensional) marginal distribution of at Note that although is a population object, we assume that it is available since we are interested in identification. Assume that the interior of the support of denoted by is not empty.
@@
As we discussed in the previous section, [dette2016testing] propose a method to test individual rationality based on the negative semidefiniteness. Their testing procedure is roughly as follows. In their Theorem 1, they show that the quadratic form of the average Slutsky matrix is identified, that is, is identified for each and If the population consists of rational individuals, then is negative semidefinite almost surely, and hence, should hold. One can reject the null hypothesis that the population is rational if an estimate of is significantly positive for some and 333To be more precise, [dette2016testing] show a stronger result that the quadratic form of the average Slutsky matrix conditional on the value of Thus, their test is more powerful than what is described here.
@@
Unfortunately, the following theorem shows that Slutsky symmetry is not testable based on the average Slutsky matrix. The proof is illustrated in the next section.
@@
Suppose satisfies Assumption B.1. There exists a -valued random demand function such that and is symmetric for all
Theorem 2.1 particularly implies that the average Slutsky matrix is not identified in our setup. The difficulty arises from the second term of (1). The expectation of the second term is written as
but it is unclear how to identify because it involves the joint distribution of demand at different income levels, and it is indeed not identifiable as we see in the proof of Theorem 2.1.
Theorem 2.1 implies that no matter what cross-sectional demand distributions satisfying the regularity condition we observe, there always exists a stochastic demand system such that it is observationally equivalent to the true demand system and its average Slutsky matrix is symmetric. Hence, there is no way to infer whether is symmetric or not from
Theorem 2.1 does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. It can be still possible that individual symmetry has observable implications through nonlinear statistics, but we leave this for future work.
@@
In addition to the marginal distribution of let us assume that the joint distribution of is available for all This setup corresponds to the situation where we can observe each individualβs choice twice. In this case, the average Slutsky matrix is identified because the RHS of (2) is identified. Hence, it is possible to test the Slutsky symmetry by testing whether is symmetric or not.
@@
Proposition 2 of [maes2024beyond] asserts that the average Slutsky matrix is not βautomaticallyβ identified. They explain that the quantity (2) is not identified in the same way that the other term is identified through the equation Their argument is not complete, as it merely rules out a particular strategy for identifying the average Slutsky matrix without addressing the possibility of alternative identification approaches. Consequently, [maes2024beyond] do not establish whether there exists an observationally equivalent demand system whose average Slutsky matrix is symmetric.
3 Construction of Stochastic Demand Systems
In this section, we illustrate how a stochastic demand system with the properties of Theorem 2.1 is constructed for the three-good case (). General cases are discussed in the appendix. Remember that for given we need to construct a -valued random element such that and is symmetric. We first focus on the first condition, the marginal compliance. Let
@@ 3 Construction of Stochastic Demand Systems
Let be such that Assumption B.1. Then, there exists a measurable function such that
1.
it is continuously differentiable in each argument,
2.
is a homeomorphism for each
3.
for each and
4.
for each
In particular, the -valued random function where complies the marginal distributions,
In the random demand function constructed in Lemma 3.1, consumersβ preference heterogeneity, or βtype,β is encoded in By the third property of consumerβs type is understood as the demand at Observe that the demand system constructed in Lemma 3.1 is degenerated in the sense that consumers are completely characterized by the demand at More specifically, if an individual demands at (s)he demands at for sure.
Although in Lemma 3.1 satisfies the marginal compliance, it does not satisfy the average Slutsky symmetry condition in general. Our strategy is to construct another random demand function satisfying both conditions by modifying The following lemma gives a condition equivalent to the symmetry.
@@
Let be such that Assumption B.1. Suppose that a -valued random function satisfies for all Then, its average Slutsky matrix is symmetric if and only if
holds.
We shall modify so that it satisfies equation (3). Fix Let be the income derivative of the demand function of consumer
where Notice that this is well-defined since is homeomorphic. The family of vector fields, combined with an initial condition, pins down the demand function, as is the unique solution to the ODE
under the global Lipschitz condition on by the Cauchy-Lipschitz theorem. Since is not equal to the RHS of (3) in general, we shall rectify with an auxiliary vector field constructed in the following lemma.
Lemma 3.3.
Fix Let be such that is smooth. There exists a vector field that is Lipschitz uniformly in and satisfies
For
solve the partial differential equation (PDE) (4) and let Consider the following ODE
Lemma 3.4.
The random function constructed in this lemma satisfies the conditions of Theorem 2.1.
In a similar idea, we can explicitly write down the identified set of the average Slutsky matrix.
Proposition 3.1.
Let be such that Assumption B.1. Suppose that a -valued random function satisfies for all The identified set of is
It is easy to see that the identified set always contains a symmetric matrix, which is consistent with the consequence of Theorem 2.1. We also observe that cross-sectional demand distributions put no restriction on the off-diagonal entries except for an trivial restriction

### v2

@@ Abstract
Slutsky symmetry and negative semidefiniteness are necessary and sufficient conditions for the rationality of demand functions. While the empirical implications of Slutsky negative semidefiniteness in repeated cross-sectional demand data are well understood, the empirical content of Slutsky symmetry remains largely unexplored. This paper takes an important first step toward addressing this gap. We show that the average Slutsky matrix is not identified and that its identified set always contains a symmetric matrix, implying that the symmetry of the average Slutsky matrix is untestable and that individual Slutsky symmetry cannot be tested through the average. Nevertheless, we demonstrate that, by imposing bounds on the income elasticity of demand, Slutsky symmetry implies a set of functional inequality constraints that are testable.
Keywords— Slutsky symmetry, rationality, cross-sectional demand, continuity equation
@@ 1 Introduction
Rationality is a central concept in economic theory, serving as a fundamental assumption in the analysis of consumer behavior. It assumes that consumers make decisions that are consistent with utility maximization, a principle that supports many economic models and empirical research. Testing this assumption is crucial for validating theoretical models and understanding consumer decision-making.
When individual demand functions are available, the Hurwicz-Uzawa theorem ([8]) provides a complete characterization of rationality. To state the theorem, let be a -dimensional price vector and be an income, both of which are relative to the price of the numeraire. Let be the -dimensional vector of quantities demanded when the price is and the income is We say that a demand function is rational if it maximizes a utility function subject to the budget constraint. According to the Hurwicz-Uzawa theorem, a demand function is rational if and only if its Slutsky matrix, which is defined as a matrix
@@
In many real-world applications, individual demand functions are not observed and only cross-sectional data is available. To investigate the population rationality, researchers have explored the rationalizability of the average demand function conditional on observable characteristics, such as price and income, treating it as if it is generated by an individual representing the population. See, for example, [13], [14], [5], and [7].
However, this approach involves some disadvantages. For instance, as Theorem 1 of [14] demonstrates, the rationality of the average demand function is irrelevant to the rationality of individual demand functions: it is possible that all individuals in a population are rational while the average demand is not, and vice versa. Moreover, it is often reported that cross-sectional mean regressions fail to explain the variation of demand adequately because of the unobserved preference heterogeneity. ([7], [6]). These facts indicate the need to investigate all the information about the heterogeneity contained in the data.
@@
Compared with testing the rationality of an individual demand function, testing whether a cross-sectional demand data is consistent with a rational population is a harder problem because individual Slutsky matrices are not observed. This difficulty raises a fundamental question: can we infer the rationality of a population from cross-sectional demand data? More specifically, is it possible to construct a statistical test with nontrivial power to detect consumer irrationality only from demand distributions?
@@
Recent research has made progress in addressing this question. [6] provide a necessary and sufficient condition for cross-sectional demand distributions to be rationalizable when only two goods are present. For cases involving more than two goods, [2] provide the empirical content of the negative semidefiniteness of the Slutsky matrix. Specifically, they show that the quadratic form of the average Slutsky matrix conditional on price and income, where is the random demand function representing the population, is identified. If an estimate of the quadratic form is significantly positive, then it means that the observed data is likely to be inconsistent with a population having an negative semidefinite Slutsky matrix—let alone rationality. These studies demonstrate the feasibility of testing rationality by means of the negative semidefiniteness of the Slutsky matrix.
However, these methods neglect the symmetry condition, leading to statistical tests that are overly conservative and fail to fully leverage the implications of rationality. The symmetry condition can be more critical to the rationality than the negative semidefiniteness, given that the former is highly sensitive to small perturbations while the latter is not. If Slutsky symmetry is testable, one should be able to construct a much more powerful test for rationality by testing both conditions.
@@
Despite its theoretical importance, the empirical content of the Slutsky symmetry condition remains poorly understood. This paper provides a first step toward elucidating the implications of individual Slutsky symmetry for cross-sectional demand distributions. The main result of this paper claims that the average Slutsky matrix is not identified, even though its quadratic form is identified, as shown in [2]. Moreover, we show that the identified set of the average Slutsky matrix always contains a symmetric matrix.
@@
An immediate consequence of our results is the fundamental untestability of individual Slutsky symmetry via the average Slutsky matrix. This fact indicates that the Slutsky symmetric condition is totally distinct to the negative semidefiniteness condition, which is testable based on the quadratic form of the average Slutsky matrix ([2]).
The untestability of the average Slutsky symmetry does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. After the first draft of this paper was made public, [4] derive the empirical content of individual Slutsky symmetry without relying on the average Slutsky matrix. They show that individual Slutsky symmetry implies a nonparametric conditional quantile restriction, which is testable.
Although the average Slutsky symmetry is untestable in general, it can become testable under additional assumptions. In particular, we consider imposing bounds on the income elasticity of demand. By restricting the patterns of demand substitution in this way, Slutsky symmetry leads to functional inequality conditions that are, in principle, testable. Constructing a formal statistical test based on these conditions is beyond the scope of this paper, but existing methods for testing functional inequalities could be applied.
A key theoretical contribution of this paper is the development of a constructive method to generate a random demand function that satisfies prescribed marginal distributions while remaining consistent with the Slutsky conditions. In general, both the levels and the derivatives of individual demand functions vary across consumers, reflecting heterogeneous preferences. The construction in this paper shows, however, that even if one fixes a deterministic rule describing how demand changes with prices and income, it is still possible to produce random demand distributions that are consistent with observed marginals and with rational behavior. The key insight is that once the local response of demand to prices and income is fixed, the distribution of demand at all other price-income pairs can be obtained from the distribution at a single reference point by following the trajectories implied by this response rule. Mathematically, this propagation can be interpreted as the solution to a system of differential equations that transports the reference distribution along price-income changes.
@@
Related literature. [8] investigate the integrability of individual demand functions and give a necessary and sufficient condition for rationality based on the Slutsky matrix. [14] considers a population that is heterogeneous in preference and provides conditions for the average demand function to be rational. [5] and [7] among others consider the estimation and inference on the average demand function under the rationality shape restriction.
@@
For cross-sectional demand distributions, [6] give a necessary and sufficient condition for observed datasets to be consistent with a rational demand system for the cases when there are only two goods. When more than two goods are present, the characterization of rationalizability is largely open. [2] provide a necessary condition that the data needs to satisfy for it to be rationalizable focusing on the negative semidefiniteness of the Slutsky matrix and propose a statistical testing for rationality. [16] construct a testing procedure of rationality based on higher order moments of demand distributions. They also observe that the average Slutsky matrix is not identified from those moments. More recently, [4] derive the empirical content of individual Slutsky symmetry without relying on the average Slutsky matrix.
@@
This paper is also related to the literature on random utility models. [18] and [17] show that the axiom of revealed stochastic preference characterizes rationalizability of stochastic choice functions defined on a finite number of choice sets. In a similar setup, [9] constructed a statistical test for the axiom. In the context of discrete choice, [1] recently gives the complete characterization of the rationalizability of demand distributions.
2 Setup and Results
2.1 Setup
@@
For we consider an economy with goods. Relative to the first good, their prices are encoded into a price vector where Let be the relative income. For given price and income a consumer demands Notice that we assume the homogeneity of demand functions at this point. We also assume Warlas’ law, i.e., the demand for the numeraire is and consequently, is assumed.
@@
In the real world, consumer’s preference heterogeneity is present. In this sense, consumer’s demand is stochastic from the perspective of an econometrician. Let be a random individual demand function drawn from a probability distribution on The econometrician is assumed to observe cross-sectional demand distributions, that is, (s)he observes the distribution of for each but no joint distribution of demands across different price-income levels is available. Regarding as a stochastic process indexed by we often call the (one-dimensional) marginal distribution of at Note that although is a population object, we assume that it is available since we are interested in identification. Assume that the interior of the support of denoted by is not empty.
@@
2.2 Main Result
As we discussed in the previous section, [2] propose a method to test individual rationality based on the negative semidefiniteness. Their testing procedure is roughly as follows. In their Theorem 1, they show that the quadratic form of the average Slutsky matrix is identified, that is, is identified for each and If the population consists of rational individuals, then is negative semidefinite almost surely, and hence, should hold. One can reject the null hypothesis that the population is rational if an estimate of is significantly positive for some and 222To be more precise, [2] show a stronger result that the quadratic form of the average Slutsky matrix conditional on the value of Thus, their test is more powerful than what is described here.
@@ 2.2 Main Result
If individual Slutsky matrix is symmetric (almost surely), so is its average. In what follows, we address the question of what we can say about the average Slutsky matrix from cross-sectional demand distributions
First, we observe that the average Slutsky matrix is not identified in an explicit way because of the second term of (1). The expectation of the second term is written as
but it is unclear how to identify because it involves the joint distribution of demand at different income levels, and it is indeed not identifiable as Theorem 2.1 implies below.
The nonidentifiability of the average Slutsky matrix does not immediately imply that we cannot say anything about its symmetry. The individual symmetry could have some empirical implications for the observable demand distributions. Unfortunately, however, there is nothing we can say about the symmetry of the average Slutsky matrix from the cross-sectional demand data. The following theorem formalizes the argument so far. The proof is given in Appendix.
@@
Let be such that Assumption B.1. The identified set of the function that maps price-income pairs to the average Slutsky matrix at is given by
where
is identified. In particular, the average Slutsky matrix is not identified at any and there necessarily exists a -valued random demand function such that and is symmetric for all
Assumption B.1 imposes regularity on the family of demand distributions. Specifically, it requires that has a smooth density, and that both the density and support vary smoothly with
@@
Proposition 2 of [16] asserts that the average Slutsky matrix is not “automatically” identified. They explain that the quantity (2) is not identified in the same way that the other term is identified through the equation Their argument is not complete, as it merely rules out a particular strategy for identifying the average Slutsky matrix without addressing the possibility of alternative identification approaches. Consequently, [16] do not establish whether there exists an observationally equivalent demand system whose average Slutsky matrix is symmetric.
2.3 Practical Implications
Testability of Slutsky symmetry. Theorem 2.1 has several useful implications. First, as stated in the theorem, the average Slutsky matrix is not identified, and the identified set is unbounded in the sense that can be arbitrarily large by is small. Moreover, Theorem 2.1 implies that no matter what cross-sectional demand distributions satisfying the regularity condition we observe, there always exists a stochastic demand system such that it is observationally equivalent to the true demand system and its average Slutsky matrix is symmetric. Consequently, there is no way to infer whether is symmetric or not from
This negative result suggests that researchers collect additional demand data in order to identify the average Slutsky matrix and test Slutsky symmetry. For example, let us assume that the joint distribution of is available for all rather than the marginal distribution This setup corresponds to the situation where analysts can observe each individual’s choice twice. In this setup, the average Slutsky matrix is identified because the RHS of (2) is identified, and hence, it is possible to test the Slutsky symmetry by testing whether is symmetric or not.
It is also worth mentioning that even though Theorem 2.1 does show that Slutsky symmetry is not testable based on the average Slutsky matrix, it does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. Indeed, individual symmetry could have observable implications through nonlinear statistics rather than the average, but we leave this for future work.
Adding income elasticity bounds. Now, we shall see that the average Slutsky symmetry can be testable by imposing bounds on the income elasticity of demand. Let
be the income elasticity of demand for th good for demand system Then, we have
Let be the income elasticity of demand It is easy to check
It is often reasonable to put bounds on the income elasticity of demand Assume that there are functions such that for all and Then we have
This inequality leads to a bound on the difference for it must be lie in the interval
Since is identified and estimable, the average Slutsky symmetry— for all and —can be tested by checking whether the null hypothesis
While constructing a statistical test for this hypothesis is beyond the scope of this paper, the literature on testing functional inequalities, such as [11], [12], and [15], will work.
3 Construction of Stochastic Demand Systems
Let be an element of the identified set of the average Slutsky matrix. To prove Theorem 2.1, it is sufficient to construct a -valued random demand function such that
for The goal of this section is to describe the construction of such a random demand function.
In overview, the construction proceeds in two steps. First, we construct a preliminary random demand function such that (M) holds but not necessarily (S). In the second step, we modify to obtain another random demand function that satisfies (S) as well as (M).
Step 1. We begin with constructing a preliminary random demand function satisfying (M). Let
Lemma 3.1.
@@ 3 Construction of Stochastic Demand Systems
In the random demand function constructed in Lemma 3.1, consumers’ preference heterogeneity, or “type,” is encoded in By the third property of consumer’s type is understood as the demand at Observe that the demand system constructed in Lemma 3.1 is degenerated in the sense that consumers are completely characterized by the demand at More specifically, if an individual demands at (s)he demands at for sure.
Before moving on to the next step, we observe that the random demand function is characterized by the solution to an ordinary differential equation (ODE). Fix Let be the income derivative of the demand function of consumer
where Notice that this is well-defined since is homeomorphic. The family of vector fields, combined with an initial condition, pins down the demand function, as the Cauchy-Lipschitz theorem implies that is the unique solution to the ODE
under the global Lipschitz condition on
Step 2. Although in Lemma 3.1 satisfies the marginal compliance (M), it does not satisfy the average Slutsky symmetry condition in general. Our strategy is to construct another random demand function satisfying both conditions by modifying To do so, we modify the ODE (3) so that its solution respects (S). More precisely, we rectify by adding an auxiliary vector field constructed in the following lemma.
@@
For let be a continuous function on Assume that is satisfied for all There exists a vector field that is Lipschitz uniformly in and satisfies
For set
which is a continuous function on under Assumption B.1. Then we have if since and
where the first equality holds by the law of motion of the ODE (3), and the second equality holds by the fact that satisfies (M). Take a vector field from Lemma 3.2 for this and let Consider a modified ODE
The following lemma shows that this flow induces a random demand function that satisfies the desired requirements.